Presentation on theme: "1 ACS Status Update STScI TIPS Meeting 19-October-2006 Ken Sembach."— Presentation transcript:
1 ACS Status Update STScI TIPS Meeting 19-October-2006 Ken Sembach
2 ACS/HRC Timeline ACS Suspended Saturday, September 23, 2006 Telemetry showed that the HRC ASPC2 board was missing its +35V input This voltage is needed for CCD bias level and reset drain operation. Loss of +35V believed to leave HRC unusable Lots of discussion about this - more on this if you want to know An ACS-free SMS was implemented the following week Thanks to scheduling team, PCs, and NICMOS Team for getting as much science as possible planned for that week ARB was convened to determine cause of suspension Failure of A2/A3 contacts to close in K1 relay Solder joints from K1 relay to ASPC2 board Problem is unrelated to June 2006 Side-1 electronics failure
3 CEB Power Switching LVPS1 LVPS2 CEB Clock1 (CCD control) Clock2 (CCD control) All boards are needed for operation of the entire detector. ASPC1 (video) ASPC2 (video) Timing (main control) Simplified – multiple diode busses in hardware MEB1 MEB2 Legend: 5V only All relays independent and switched by individual RIU commands
4 K1 relay ASPC1 (A,B half of CCD) ASPC2 (C,D half of CCD)
5 Relay Operation Rocker Assembly Pivot point Glass Beads: Used to push on movable contact A C B D Pushes +35V Pole Pushes Telemetry Pole Contact made by spring action of movable center arm to relaxed state Contact made by pushing action of rocker arm Shown in LVPS2 Selection Position (X coil commanded last) A3 A1 A2D2 D1 D3 (LVPS2 +35V) (LVPS1 +35V)
6 ACS/HRC Timeline WFC returned to service on October 1, 2006 Not affected by HRC relay issue HRC returned to service on October 15, 2006 Single toggle of relay restored electrical contact ARB has recommended never toggling problematic relay again SBC is functional but powered off ARB is examining options for bringing SBC on line Previous use of SBC required toggle of problematic HRC relay to allow HRC to be turned off for SBC observations Power / current / thermal issues are being worked Option 1: CCDs and SBC on simultaneously Option 2: CCDs off when SBC is on
7 ACS/HRC Post Recovery Performance Internal bias and dark exposures from ACS Daily Monitor Program (10758) have been examined Data were obtained after the ACS anneal on October 16, 2006 Data were read out through the C amplifier Both GAIN=2 and GAIN=4 data were obtained. Readnoise and dark current track historical averages Bias levels are slightly higher than historical average, but this is expected Bias level typically higher for a short time (few hours) after anneals Bias level is expected to drop back to historical average by next set of observations No impact on science data or performance Data look perfectly normal, with no anomalies noted (Marco Sirianni)
8 ACS/HRC Bias Images (16-Oct-2006) The two bias images above were obtained after resumption of HRC operations. Both were obtained with GAIN= 2 after the ACS anneal. These bias images look perfectly normal. The data were read out through the C amplifier.
9 ACS/HRC Readnoise and Bias Level (GAIN=2) The above plots show the readnoise (left) and bias level (right) for historical Side-2 data [blue points] and post-recovery data [yellow points]. The solid red lines are the historical averages, and the dotted red lines indicate the historical ±3 sigma excursions.
10 ACS/HRC Readnoise and Bias Level (GAIN=4) The above plots show the readnoise (left) and bias level (right) for historical Side-2 data [blue points] and post-recovery data [yellow points]. The solid red lines are the historical averages, and the dotted red lines indicate the historical ±3 sigma excursions.
11 ACS/HRC Post Recovery Performance First post-recovery science data were obtained October 18, 2006 G800L observations of SN1987A (Kirshner SAINTs program; GO10867) Historical data were pulled for comparison Exposures dithered to remove cosmic rays Data look perfectly normal, with no anomalies noted (Ron Gilliland)